I need to put 8x+5y=30 in to y=
please help i have a big test tomorrow a really need help so i can study from my notes. thanks

Answer by bennettc18(2)

Move x over to the other side making it 5y=8x+30
Next divide the entire problem by 5 so the result would be 5/5y=8/5X+30/5
After pluging those in you get y=8/5x+6